OpenStudy (anonymous):

Find the area of a regular hexagon with the given measurement. 2 sqrt 3 apothem A =

OpenStudy (jdoe0001):

all you're given is the apothem, right? nothing else

OpenStudy (anonymous):

I'll give you the formula of the hexagon which is \[A=\frac{ 3\sqrt{3} }{ 2 }a ^{2}\]

OpenStudy (anonymous):

@jdoe0001 yep thats it

OpenStudy (jdoe0001):

ok so...let's see

OpenStudy (jdoe0001):

\(\bf \textit{area of a regular polygon}=nr^2\cdot tan\left(\cfrac{180}{n}\right)\quad \begin{cases} n=\textit{number of sides}\\ r=apothem \end{cases}\)

OpenStudy (jdoe0001):

is a HEXAgon.... so it has 6 sides and you have the apothem plug and chug

OpenStudy (anonymous):

is it 30 sqrt 3?

OpenStudy (jdoe0001):

well... keep in mind that is a bit more than 30 \(\bf \textit{area of a regular polygon}=nr^2\cdot tan\left(\cfrac{180}{n}\right)\quad \begin{cases} n=\textit{number of sides}\to 6\\ r=apothem\to 2\sqrt{3} \end{cases} \\ \quad \\ 6\cdot (2\sqrt{3})^2\cdot tan\left(\cfrac{180}{6}\right)\implies 6\cdot 2^2(\sqrt{3})^2\cdot tan\left(\cfrac{\cancel{ 180 }}{\cancel{ 6 }}\right) \\ \quad \\ 6\cdot 4\cdot 3\cdot tan\left(30\right)\)

OpenStudy (jdoe0001):

anyhow.. need to dash...but is pretty straightfoward from there

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

that one is tricky